OpenStudy (anonymous):

PLEASE help me ? A bird sees a worm on the ground 50 m. below. The bird dives for the worm at a speed of 6 m/sec. The function h(t)= - 4.9t^2+6t+50, where h(t) is the height of the bird at time t. How long does the worm have before he is eaten?

OpenStudy (ybarrap):

Set h(t) = 0 and solve for t.

OpenStudy (anonymous):

how would i write that with this problem?

OpenStudy (ybarrap):

note that h(0) = 50 meters. That is, at the start of the fall, the bird is 50 meters.

OpenStudy (ybarrap):

h(t)=0=-4.9t^2+6t+50

OpenStudy (ybarrap):

-4.9t^2+6t=-50

OpenStudy (ybarrap):

-4.9t^2+6t + (6/(2*4.9)^2=-50+(6/(2*-4.9)^2

OpenStudy (anonymous):

-18.01 fo the first part of the problem

OpenStudy (anonymous):

is that right

OpenStudy (ybarrap):

Bottom, line. You'll have a quadratic, one positive, one negative. Take the positive: Solve for t over the real numbers: 0 = -4.9 t^2+6 t+50 Reverse the equality in 0 = -4.9 t^2+6 t+50 in order to isolate t to the left hand side. 0 = -4.9 t^2+6 t+50 is equivalent to -4.9 t^2+6 t+50 = 0: -4.9 t^2+6 t+50 = 0 Write the quadratic equation in standard form. Divide both sides by -4.9: t^2-1.22449 t-10.2041 = 0. Solve the quadratic equation by completing the square. Add 10.2041 to both sides: t^2-1.22449 t = 10.2041 Take one half of the coefficient of t and square it, then add it to both sides. Add 0.374844 to both sides: t^2-1.22449 t+0.374844 = 10.5789 Factor the left hand side. Write the left hand side as a square: (t-0.612245)^2 = 10.5789 Eliminate the exponent on the left hand side. Take the square root of both sides: t-0.612245 = 3.25253 or t-0.612245 = -3.25253 Look at the first equation: Solve for t. Add 0.612245 to both sides: t = 3.86477 or t-0.612245 = -3.25253 Look at the second equation: Solve for t. Add 0.612245 to both sides: Answer: | | t = 3.86477 or t = -2.64028

OpenStudy (anonymous):

wow I had to do all that ????????????? and which one is the answer how can it be 2 ?

OpenStudy (ybarrap):

Answer: | | t = 3.86477 or t = -2.64028

OpenStudy (ybarrap):

Take the positive.

OpenStudy (ybarrap):

Note that -4.9t^2+6t+50 at t = 3.86477 is zero. That is, the bird is on the ground and h(3.86477)=0.

OpenStudy (anonymous):

so it can be 0

OpenStudy (ybarrap):

This shows it.

OpenStudy (anonymous):

why don't you use the quadratic formula....

OpenStudy (anonymous):

what is it

OpenStudy (ybarrap):

Here's a plot of the function for t=0 to t=4:

OpenStudy (ybarrap):

Note that at the beginning of time it is 50 meters up and at 3.9 seconds, it hit the ground (i.e. h(t) = 0)

OpenStudy (anonymous):

I'll try 3.9...

OpenStudy (ybarrap):

If you learn the quadratic formulat as @paigeRG indicates, the solution is so much easier to solve.

OpenStudy (anonymous):

idk if i should round it to 4.

OpenStudy (ybarrap):

Then the bird will go slightly past the ground (using 4).

OpenStudy (ybarrap):

That would probably hurt the bird.

OpenStudy (anonymous):

lol true..

OpenStudy (anonymous):

http://www.purplemath.com/modules/quadform.htm <-------- heres a website that explains the quadratic formula, just so you know for future reference.... if you don't want to use it now

OpenStudy (anonymous):

Dude... @ybarrap you put your website before mine, not cool

OpenStudy (anonymous):

ok how would i plug my numbers in for the equation ? whats a,b,c???

OpenStudy (ybarrap):

If it's any consolation, I think your source is a lot friendlier. @xoxox123, use @paigeRG 's link first.

OpenStudy (ybarrap):

I'll let @paigeRG take over from here :)

OpenStudy (anonymous):

Well ill be here if you need anything @xoxox123

OpenStudy (anonymous):

ok good cause i need your help. can u help me plug the numbers from my equation to the formula

OpenStudy (anonymous):

@paigeRG

OpenStudy (anonymous):

yes, so in the equation, there are three letters a b and c... a is the first term of the equation b is the second and c is the third term so A=-4.9 B=6 and C=50

OpenStudy (anonymous):

okay i understand that..

OpenStudy (anonymous):

\[-b \pm \sqrt{b^2-4(a)(c)}\] and so that would be over \[2(a)\]

OpenStudy (anonymous):

now you would plug in the terms... and since there is a negative times a negative it would turn possative -4(-4.9)(50) -----> so it would be 36+980

OpenStudy (anonymous):

to get what ..?

OpenStudy (anonymous):

1016?

OpenStudy (anonymous):

are you there ??

OpenStudy (anonymous):

whats that sign next to -b ?

OpenStudy (anonymous):

sorry I had to eat dinner

OpenStudy (anonymous):

1016 is correct so now what you would do is find the two answers, that's what the sign nest to the -6 is. it stands for plus of minus.

OpenStudy (anonymous):

itd okay and the answer is going to be 2 ?

OpenStudy (anonymous):

\[-6\pm \sqrt{1016 }\] over 2(-4.9)=-9.8

OpenStudy (anonymous):

That plus of mius sign i never seen it before.... how do you put it in the calc

OpenStudy (anonymous):

so now what you would do is do \[-6+\sqrt{1016} \div -9.8\] thus getting your first x intercept, and for your second intercept you would do \[-6-\sqrt{1016}\div -9.8\]

OpenStudy (anonymous):

It say error when i put it in the calc ??

OpenStudy (anonymous):

so your two answers are -2.6 and 3.9 (both rounded to nearest tenth)

OpenStudy (anonymous):

wait nvm i got the first one i got -9.25252601

OpenStudy (anonymous):

you cant really put it into a calculator its just one of those formulas you have to memorize... our teacher made us memorize it with a song...

OpenStudy (anonymous):

all you do is plug the terms into the correct places and basically simplify the equation down to where you can figure the points, if you haven't learned it yet you will soon so no worries

OpenStudy (anonymous):

well how did you solve -6+√1016/-9.8

OpenStudy (anonymous):

well that part you can solve with a calculator

OpenStudy (anonymous):

thats what i did and got that longg decimal

OpenStudy (anonymous):

that's fine I got long decimals too, but I rounded them to the nearest tenth... I told you the answers you should have gotten did you see them?

OpenStudy (anonymous):

yeah i saw them i still have 3 more problems then im a submit my quiz. and see if its right

OpenStudy (anonymous):

I still have 2 more problems dealing with the same thing you just did //:

OpenStudy (anonymous):

oh goodness

OpenStudy (anonymous):

i knoww >.< can u please help me ??

OpenStudy (anonymous):

it would be really helpfull :(

OpenStudy (anonymous):

I can try but I have math stuff to work on too so ill see what I can do

OpenStudy (anonymous):

OK thanks so much here it is.... A rocket is fired straight up from a 60 ft. platform with an initial velocity of 96 ft/sec. The height of the rocket, h(t), is found using the function h(t) = - 16t^2 + 96t + 60 where t is the time in seconds. Find the number of seconds required to reach the maximum height.

OpenStudy (anonymous):

Maximum height... I always hated finding the height of stuff... Ill see what I can do...

OpenStudy (anonymous):

i hate it too but okay ! thanks

OpenStudy (anonymous):

okay I found the x intercepts im just trying to figure out the height

OpenStudy (anonymous):

if you dont know that one can u help with this one ??? A ball is dropped from the top of a 550 ft. building. The function h(t ) = - 16t^2 + 50 models the height of the ball, h(t) (in feet), at any given time, t (in seconds). How long does it take the ball to reach the ground? Round your answer to the nearest hundredth of a second.

OpenStudy (anonymous):

do you have multiple choices

OpenStudy (anonymous):

no sadly its written response

OpenStudy (anonymous):

for the rocket one Im positive its 3

OpenStudy (anonymous):

wait!

OpenStudy (anonymous):

okay I will try it

OpenStudy (anonymous):

noooo

OpenStudy (anonymous):

what ??

OpenStudy (anonymous):

im triple checking, just wait

OpenStudy (anonymous):

yup its 3

OpenStudy (anonymous):

okay time for the last one

OpenStudy (anonymous):

okay I hope its right :)

OpenStudy (anonymous):

okay question, is the equation for the last question correct? is it supposed to be -16t^2+50 or is it supposed to be -16t^2+550.... is it plus 50 or plus 550

OpenStudy (anonymous):

its 50

OpenStudy (anonymous):

your positive? because it says its dropped from a 550 ft building... unlesing is 50 ft tall and not 550

OpenStudy (anonymous):

oh wait i thought your talking about the equation was 50 ? but yess the ball dropped from a 500 ft building.. sorry misread

OpenStudy (anonymous):

500 ft... okay

OpenStudy (anonymous):

Are you possssssiiitivvve the equation has 50 at the end...

OpenStudy (anonymous):

NO! its 550!! omg im going to rewrite this for you.

OpenStudy (anonymous):

well do you have the link> just copy and paste it so I can see for myself what it says

OpenStudy (anonymous):

I cant because you would need the password for my class. I'll copy it exactly the way it is. A ball is dropped from the top of a 550 ft. building. The function h(t)= -16t^2+50 models the height of the ball, h(t) (in feet), at any given time, t (in seconds). How long does it take the ball to reach the ground? Round your answer to the nearest hundredth of a second.

OpenStudy (anonymous):

1.77

OpenStudy (anonymous):

is it rounded.

OpenStudy (anonymous):

Yes its rounded....I still think the equation is wrong but what ever... hope you get it right

OpenStudy (anonymous):

I just submited it

OpenStudy (anonymous):

And did I help you in anyway shape or form

OpenStudy (anonymous):

3 is right , but 1.77 is not when i got it wrong it said to Graph the function. Where on the graph is the ground represented?

OpenStudy (anonymous):

x axis

OpenStudy (anonymous):

I graphed it on a graphing calculator and checked it a dozen times, but it still sounded wrong... the equation sounded wrong I mean

OpenStudy (anonymous):

no the problem is right i made sure i wrote it right.. and i got the bird one wrong!!!